A consumer upset with the latest trend of postal r
Answer #1

We know that the formula for the Buoyancy force is given by

Fb =ma+mg

pgV =ma+mg

pgV =m(a+g)

Then mass (m) =pgV/(a+g) =(1.16kg/m3)(9.81m/s2)(6*10-2m3)/(2.65m/s2+9.81m/s2)=0.054797kg

Now the mass of the air balloon is given by

m' =m -mp =0.054797kg -0.04kg=0.014797kg

Then the density of the gas consumer is (p) =mass/volume =0.014797kg/6*10-2m3=0.2466kg/m3